OpenStudy (anonymous):

Solve. x2 + 5x = –6 x = 5 and x = 1 x = –5 and x = –1 x = 3 and x = 2 x = –3 and x = –2

OpenStudy (anonymous):

first you want the equation ot equal zero so add six to both sides to get x^2+5x+6 = 0 now you factor to get (x+2)(x+3)=0 now because the factors multiply to zero, you know that one of them must equal zero so you write x+2 = 0 and x+3=0 x=-2 and x=-3